# Towards a theory of model distillation

### v1

@@ 5.4.1 Sample complexity of learning is determined by the VC dimension
The VC dimension gives both upper and lower bounds for the number of samples needed to (agnostically) learn [96, 22]. We follow the presentation of [1].
Proposition 5.11 (VC dimension controls sample complexity of learning; Theorem 5.6 in [1]).
There are constants $c_{1},c_{2}>0$ such that for any $80<\epsilon<1/$ 8 and $1000<\delta<1/$ 100, the minimum number of samples $n(\epsilon,\delta)$ needed to $(\epsilon,\delta)$-learn concept class ${\mathcal{G}}$ with hypothesis class ${\mathcal{G}}$ is: $\frac{c_{1}}{\epsilon}\left(\mathrm{VCdim}({\mathcal{G}})+\ln(1/\delta)\right)\leq n(\epsilon,\delta)\leq\frac{c_{2}}{\epsilon}\left(\mathrm{VCdim}({\mathcal{G}})\ln(1/\epsilon)+\ln(1/\delta)\right)\,.$ If $\mathrm{VCdim}({\mathcal{G}})=\infty$, then ${\mathcal{G}}$ is not learnable in finitely many samples.
@@
5.4.2 Is there an analogue for distillation?
Is there a quantity analogous to the VC dimension that determines how many samples are needed to distill? This question can be asked in both the standard and agnostic settings.
Open problem 1.
Find a combinatorial property of the pair $({\mathcal{F}}$, ${\mathcal{G}})$ that controls the sample complexity of distilling ${\mathcal{F}}$ into ${\mathcal{G}}$.
Open problem 2.
Find a combinatorial property of the pair $({\mathcal{F}}$, ${\mathcal{G}})$ that controls the sample complexity of agnostically distilling ${\mathcal{F}}$ into ${\mathcal{G}}$.
In the case that the source class can be perfectly distilled into the target class (Definition 5.3), then we have resolved the open problems in Section 5.2. For finite input spaces we have proved that ${\mathcal{F}}$ is distillable into ${\mathcal{G}}$ in 0 samples, and for countable input spaces we have proved a $O(\log(1/\delta)/\epsilon)$-sample upper bound, with a matching lower bound. The question is open when perfect distillation is not possible. We make partial progress below.
5.4.3 The VC dimension of the Pareto frontier does not characterize distillation
We give some preliminary results for open problems 1 and 2, but do not fully resolve these problems. In this section, we show that the sample complexity of agnostic distillation is bounded above by a natural combinatorial quantity called the “VC dimension of the Pareto frontier”, which can be much smaller than the VC dimension. However, we prove that there is no corresponding lower bound based on the VC dimension of the Pareto frontier.
First, in Lemma 5.13, we reduce distillation without loss of generality to the case where the label alphabet is binary and the source class is a singleton set. [^7] The reduction breaks when distillation is extended to different loss functions as discussed in Section 6. The relevant notation for our reduction is as follows. For a function $f\in{\mathcal{Y}}^{{\mathcal{X}}}$ and a class of functions ${\mathcal{G}}\subseteq{\mathcal{Y}}^{{\mathcal{X}}}$, define the binary-valued class of functions that represents the error pattern of the hypotheses with respect to $f\in{\mathcal{F}}$:
@@ 5.4.3 The VC dimension of the Pareto frontier does not characterize distillation
Under this lens, one can see that some target functions have strictly more errors than others with respect to the source concept. Thus, an optimal distillation algorithm will always output a target function whose error pattern does not dominate any of the other target functions. Formally, we can reduce to the Pareto frontier of the set of target functions:
@@
Let the source class be =${\mathcal{F}}=\{\mathrm{zero}\}\subseteq\{0,1\}^{{\mathcal{X}}}$ and let the target class be ${\mathcal{G}}\subseteq\{0,1\}^{{\mathcal{X}}}$. We define the Pareto partial order $\prec_{Par}$ between pairs of distinct functions $g,g^{\prime}\in{\mathcal{G}}$. We say that $g\prec_{Par}g^{\prime}$ if whenever $g(x)=0$ then also $g^{\prime}(x)=0$. The Pareto frontier of the class ${\mathcal{G}}$ is defined as the set of maximal functions: $\mathrm{PF}({\mathcal{G}})=\{g\in{\mathcal{G}}\mid\mbox{ there isno }g^{\prime}\in{\mathcal{G}}\mbox{ s.t. }g\prec_{Par}g^{\prime}\}\,.$ Then the following are equivalent: (a) Class =${\mathcal{F}}=\{\mathrm{zero}\}$ is $(\epsilon,\delta)$-distillable into class ${\mathcal{G}}$ in $n$ samples. (b) Class =${\mathcal{F}}=\{\mathrm{zero}\}$ is $(\epsilon,\delta)$-distillable into class $\mathrm{PF}({\mathcal{G}})$ in $n$ samples. Furthermore, the same is true for agnostic distillation instead of distillation.
@@
(a) $\implies$ (b): by definition for any $g\in{\mathcal{G}}$ there is $g^{\prime}\in\mathrm{PF}({\mathcal{G}})$ such that for all distributions ${\mathcal{D}}$ we have $\mathrm{error}_{{\mathcal{D}}}(g^{\prime})\leq\mathrm{error}_{{\mathcal{D}}}(g)$. Thus an algorithm distilling ${\mathcal{F}}$ with ${\mathcal{G}}$ can be converted to outputting a corresponding function in $\mathrm{PF}({\mathcal{G}})$ instead without lowering the error.
(b) $\implies$ (a): since $\mathrm{PF}({\mathcal{G}})\subseteq{\mathcal{G}}$. For agnostic distillation, note the optimum is in $\mathrm{PF}({\mathcal{G}})$. ∎
The above reductions allow us to upper-bound the sample complexity of agnostic distillation in terms of a new combinatorial quantity: the VC dimension of the Pareto frontier. [^8] We remark that Theorem 5.15 does not apply to non-agnostic distillation, which may be impossible in some cases. The example in Observation 5.7 demonstrates that distilling may be impossible even when $\mathrm{VCdimPF}({\mathcal{F}};{\mathcal{G}})=1$.
@@
For any source class ${\mathcal{F}}\subseteq{\mathcal{Y}}^{{\mathcal{X}}}$ and any target class ${\mathcal{G}}\subseteq{\mathcal{Y}}^{{\mathcal{X}}}$, define VCdimPF(F;G) = max_f F VCdim(PF(f G)) . Then there is a constant $C>0$ such that for any $400<\epsilon<1/$ 40 and $200<\delta<1/$ 20, it is possible to agnostically $(\epsilon,\delta)$-distill class ${\mathcal{F}}$ into class ${\mathcal{G}}$ in $n(\epsilon,\delta)$ samples, where n(ϵ,δ) Cϵ2 (VCdimPF(F;G) + ln(1/δ)) .
@@
By Proposition 5.12, for each $f\in{\mathcal{G}}$ the class $\mathrm{PF}(f\oplus{\mathcal{G}})$ can be agnostically $(\epsilon,\delta)$-learned with $n(\epsilon,\delta)$ samples. By Theorem 5.1, agnostic distillation reduces to agnostic learning so $\{\mathrm{zero}\}$ can be agnostically $(\epsilon,\delta)$-distilled into $f\oplus{\mathcal{G}}$ with $n(\epsilon,\delta)$ samples. This concludes the proof by the reductions in Lemmas 5.13 and 5.14. ∎
@@
  • In the example of Theorem 5.2 for when agnostic distillation can be done with 0 samples, we have equal source and target classes ${\mathcal{F}}={\mathcal{G}}$. In this case notice that $\mathrm{VCdimPF}({\mathcal{F}};{\mathcal{G}})=\max_{f\in{\mathcal{F}}}\mathrm{VCdim}(\mathrm{PF}(f\oplus{\mathcal{G}}))=\max_{f\in{\mathcal{F}}}\mathrm{VCdim}(\{\mathrm{zero}\})=0$.
  • In the example of Theorem 5.8 for when agnostic distillation may require a high number of samples, =${\mathcal{F}}=\{\mathrm{zero}\}$, and ${\mathcal{G}}$ is large and all its functions are on the Pareto frontier. Thus, $\mathrm{VCdimPF}({\mathcal{F}};{\mathcal{G}})=\mathrm{VCdim}(\mathrm{PF}({\mathcal{G}}))=\mathrm{VCdim}({\mathcal{G}})=m$, which is large.
@@
It can be seen that $g_{S,T,U}$ Pareto-dominates $g_{S^{\prime},T^{\prime},U^{\prime}}$ if and only if $S\subseteq S^{\prime}$, $T\subseteq T^{\prime}$, and $U\subseteq U^{\prime}$ and this containment is strict for one of $S,T,U$. All of the functions in ${\mathcal{G}}$ are Pareto-incomparable, so $\mathrm{PF}({\mathcal{G}})={\mathcal{G}}$.
Thus $\mathrm{VCdimPF}({\mathcal{F}};{\mathcal{G}})=\mathrm{VCdim}(\mathrm{PF}({\mathcal{G}}))=\mathrm{VCdim}({\mathcal{G}})\geq\mathrm{VCdim}({\mathcal{G}}_{1})=\infty$ because for any $m$ the set $\{(1,1),(1,2),\ldots,(1,m)\}$ is shattered by ${\mathcal{G}}_{1}$.
@@
To summarize, we have shown that, for statistical purposes, distillation reduces to considering the Pareto frontier and also to the case =${\mathcal{F}}=\{\mathrm{zero}\}$ and ${\mathcal{Y}}=\{0,1\}$. We have also shown sample complexity upper bounds for agnostic distillation in terms of a new combinatorial quantity $\mathrm{VCdimPF}$. However, we have also shown that $\mathrm{VCdimPF}$ fails to characterize the sample complexity of distillation. Thus, we leave problems 1 and 2 on fully characterizing the sample complexity of distillation open for future work.
@@ 6 Extensions and future directions
We hope that this paper encourages further theoretical and empirical exploration into distillation to help realize its promise of improved efficiency and trustworthiness in deep learning. Below, we outline some of the many open directions for future work.
@@
  • Combinatorial characterization of sample complexity. An analogue of VC-dimension for (agnostic) distillation remains open (see open problems 1 and 2).
@@ Acknowledgements
I am very grateful to Emmanuel Abbe, Kiril Bangachev, Guy Bresler, Melihcan Erol, Nati Srebro for helpful research discussions. I am also thankful to Guy Bresler for feedback on the exposition. This work was supported by an NSF Graduate Research Fellowship under NSF grant 1745302.

### v2

@@ 5.3 Bounding sample complexity of agnostic distillation using Pareto frontier
The VC dimension gives both upper and lower bounds for the number of samples needed to agnostically learn [101, 23]. We follow the presentation of [1].
@@
In this section, we show that the sample complexity of distillation is bounded above by a natural combinatorial quantity called the “VC dimension of the Pareto frontier”, which can be much smaller than the VC dimension that controls the sample complexity of learning. However, we prove that this is not tight – there is no corresponding lower bound based on the VC dimension of the Pareto frontier.
First, in Lemma 5.11, we reduce distillation without loss of generality to the case where the label alphabet is binary and the source class is a singleton set. [^7] The reduction breaks outside of the classification setting, as discussed in Section 6. The relevant notation for our reduction is as follows. For a function $f\in{\mathcal{Y}}^{{\mathcal{X}}}$ and a class of functions ${\mathcal{G}}\subseteq{\mathcal{Y}}^{{\mathcal{X}}}$, define the binary-valued class of functions that represents the error pattern of the hypotheses with respect to $f\in{\mathcal{F}}$:
@@
Under this lens, one can see that some target functions have strictly more errors than others with respect to the source concept. Thus, an optimal distillation algorithm will always output a target function whose error pattern does not dominate any of the other target functions. Formally, we can reduce to learning with any Pareto-dominating set of target functions:
Definition 5.12.
Let the source class be =${\mathcal{F}}=\{\mathrm{zero}\}\subseteq\{0,1\}^{{\mathcal{X}}}$ and let the target class be ${\mathcal{G}}\subseteq\{0,1\}^{{\mathcal{X}}}$. We define the Pareto partial order $\preceq_{Par}$ between pairs of functions $g,h\in{\mathcal{G}}$. We say that $g\preceq_{Par}h$ if whenever $g(x)=0$ then also $h(x)=0$. A subset ${\mathcal{H}}\subseteq{\mathcal{G}}$ is said to Pareto-dominate ${\mathcal{G}}$ if for all $g\in{\mathcal{G}}$ there is $h\in{\mathcal{H}}$ such that $g\preceq_{Par}h$.
@@
For any Pareto-dominating subset ${\mathcal{H}}\subseteq{\mathcal{G}}$, the following are equivalent: (a) Class =${\mathcal{F}}=\{\mathrm{zero}\}$ is $(\epsilon,\delta)$-distillable into class ${\mathcal{G}}$ in $n$ samples. (b) Class =${\mathcal{F}}=\{\mathrm{zero}\}$ is $(\epsilon,\delta)$-distillable into class ${\mathcal{H}}$ in $n$ samples. Furthermore, the same is true for agnostic distillation instead of distillation.
@@
(a) $\implies$ (b): by definition for any $g\in{\mathcal{G}}$ there is $h\in{\mathcal{H}}$ such that for all distributions ${\mathcal{D}}$ we have $\mathrm{error}_{{\mathcal{D}}}(h)\leq\mathrm{error}_{{\mathcal{D}}}(g)$. Thus an algorithm distilling ${\mathcal{F}}$ with ${\mathcal{G}}$ can be converted to outputting a corresponding function in ${\mathcal{H}}$ instead without lowering the error.
(b) $\implies$ (a): since ${\mathcal{H}}\subseteq{\mathcal{G}}$. For agnostic distillation, note the optimum is in ${\mathcal{H}}$. ∎
The above reductions allow us to upper-bound the sample complexity of agnostic distillation in terms of a new combinatorial quantity: the VC dimension of the Pareto frontier. [^8] We remark that Theorem 5.14 does not apply to non-agnostic distillation, which may be impossible in some cases. The example in Observation 5.1 demonstrates that distilling may be impossible even when $\mathrm{VCdimPF}({\mathcal{F}};{\mathcal{G}})=1$.
@@
For any source class ${\mathcal{F}}\subseteq{\mathcal{Y}}^{{\mathcal{X}}}$ and any target class ${\mathcal{G}}\subseteq{\mathcal{Y}}^{{\mathcal{X}}}$, define $\mathrm{VCdimPF}({\mathcal{F}};{\mathcal{G}})=\max_{f\in{\mathcal{F}}}\inf_{{\mathcal{H}}}\mathrm{VCdim}(f\oplus{\mathcal{H}})\,,$ where the infinum is over ${\mathcal{H}}\subseteq{\mathcal{G}}$ such that $f\oplus{\mathcal{H}}$ Pareto-dominates $f\oplus{\mathcal{G}}$. Then there is a constant $C>0$ such that for any $400<\epsilon<1/$ 40 and $200<\delta<1/$ 20, it is possible to agnostically $(\epsilon,\delta)$-distill class ${\mathcal{F}}$ into class ${\mathcal{G}}$ in $n(\epsilon,\delta)$ samples, where $n(\epsilon,\delta)\leq\frac{C}{\epsilon^{2}}(\mathrm{VCdimPF}({\mathcal{F}};{\mathcal{G}})+\ln(1/\delta))\,.$
@@
If $\mathrm{VCdimPF}({\mathcal{F}};{\mathcal{G}})$ is infinite, then the bound is vacuous and there is nothing to prove. If not, for each $f\in{\mathcal{F}}$, there is ${\mathcal{H}}_{f}\subseteq{\mathcal{G}}$ such that $\mathrm{VCdim}(f\oplus{\mathcal{H}}_{f})\leq\mathrm{VCdimPF}({\mathcal{F}};{\mathcal{G}})$ and $f\oplus{\mathcal{H}}_{f}$ Pareto-dominates $f\oplus{\mathcal{G}}$. By Proposition 5.10, for each $f\in{\mathcal{G}}$ the class $f\oplus{\mathcal{H}}_{f}$ can be agnostically $(\epsilon,\delta)$-learned with $n(\epsilon,\delta)$ samples. By ignoring the label information, agnostic distillation reduces to agnostic learning so $\{\mathrm{zero}\}$ can be agnostically $(\epsilon,\delta)$-distilled into $f\oplus{\mathcal{H}}_{f}$ with $n(\epsilon,\delta)$ samples. Combining this with the reductions in Lemmas 5.11 and 5.13 concludes the proof. ∎
@@
  • When we have equal source and target classes ${\mathcal{F}}={\mathcal{G}}$, distillation can be done in 0 samples as we saw in Theorem 5.4. In this case notice that $\mathrm{VCdimPF}({\mathcal{F}};{\mathcal{G}})\leq\max_{f\in{\mathcal{F}}}\inf_{{\mathcal{H}}}\mathrm{VCdim}(f\oplus{\mathcal{H}})=\max_{f\in{\mathcal{F}}}\mathrm{VCdim}(\{\mathrm{zero}\})=0$.
  • In the example of Theorem 5.6 for when agnostic distillation may require a high number of samples, =${\mathcal{F}}=\{\mathrm{zero}\}$, and ${\mathcal{G}}$ is large and all its functions are on the Pareto frontier – i.e., there is no proper subset ${\mathcal{H}}\subseteq{\mathcal{G}}$ such that ${\mathcal{H}}$ Pareto-dominates ${\mathcal{G}}$. Thus, $\mathrm{VCdimPF}({\mathcal{F}};{\mathcal{G}})=\mathrm{VCdim}({\mathcal{G}})=m$, which is large.
@@
It can be seen that $g_{S,T,U}$ Pareto-dominates $g_{S^{\prime},T^{\prime},U^{\prime}}$ if and only if $S\subseteq S^{\prime}$, $T\subseteq T^{\prime}$, and $U\subseteq U^{\prime}$ and this containment is strict for one of $S,T,U$. All of the functions in ${\mathcal{G}}$ are Pareto-incomparable, so there is no proper subset of ${\mathcal{G}}$ that Pareto-dominates ${\mathcal{G}}$.
Thus $\mathrm{VCdimPF}({\mathcal{F}};{\mathcal{G}})=\mathrm{VCdim}({\mathcal{G}})\geq\mathrm{VCdim}({\mathcal{G}}_{1})=\infty$, since the $\{(1,1),(1,2),\ldots,(1,m)\}$ is shattered by ${\mathcal{G}}_{1}$ for all $m$.
@@
To summarize, we have shown that, for statistical purposes, distillation reduces to considering the Pareto-dominating subsets of ${\mathcal{G}}$ and also to the case =${\mathcal{F}}=\{\mathrm{zero}\}$ and ${\mathcal{Y}}=\{0,1\}$. We have also shown sample complexity upper bounds for agnostic distillation in terms of a new combinatorial quantity $\mathrm{VCdimPF}$. However, we have also shown that $\mathrm{VCdimPF}$ fails to characterize the sample complexity of distillation.
5.4 Impossibility of characterizing sample complexity of agnostic distillation
Is there a simple combinatorial property that characterizes the sample complexity of agnostic distillation? It turns out that the answer is no, in a certain technical sense. In Lemma 5.11, we have reduced to the case where =${\mathcal{F}}=\{\mathrm{zero}\}$ and ${\mathcal{Y}}=\{0,1\}$. Thus, understanding the sample complexity of agnostic distillation reduces to understanding when, given $m$ i.i.d. samples from ${\mathcal{D}}$, we can find a hypothesis $g\in{\mathcal{G}}$ such that with probability at least $1-\delta$, we have
$\mathbb{E}_{x\sim{\mathcal{D}}}[g(x)]-\max_{g^{*}\in{\mathcal{G}}}\mathbb{E}_{x\sim{\mathcal{D}}}[g^{*}(x)]\leq\epsilon\,.$
In personal communication, Fan Chen and Sasha Rakhlin [38] have observed that this is equivalent to the “Expectation Maximization Problem” (EMX) studied in [28], which in turn is equivalent to learning in the “General Learning Setting” of Vapnik [99, 92]. By [28, Theorem 8], this implies that there is no “finite character property” (given by a first order formula provable by ZFC) analogous to the VC-dimension that characterizes the sample complexity of agnostic distillation.
@@ 6 Extensions and future directions
We hope that this work encourages further theoretical and empirical exploration into distillation. Below, we outline some of the many open directions for future work.
@@ Acknowledgements
I am very grateful to Fan Chen and Sasha Rakhlin, who pointed out the connection between the statistical complexity of distillation and the statistical complexity in the General Learning Setting of Vapnik [38]. I am also very grateful to Emmanuel Abbe, Kiril Bangachev, Guy Bresler, Melihcan Erol, and Nati Srebro for helpful research discussions. I am also thankful to Guy Bresler for feedback on the exposition. This work was supported by an NSF Graduate Research Fellowship under NSF grant 1745302.